1.You want to buy XYZ public company.  The company currently has a dividend of $3.12, but has indicated they will increase the dividend by 3.1% each year.  You expect an investment return of 9%.  How much is this stock worth?

2. You are looking to issue stock in your company.  Your company is a private school organization.  In order to get an initial price you gathered the PE ratios of five different public companies that do the same thing as you.  ABC School with a PE of 7.3, XYZ Better School with a PE Ratio of 8.6, Top Grades with a PE of 9.1, You pay we Grade, with a PE ratio of 7.1, F is the Best, with a PE Ratio of 5.9 and Way above you Means, with a PE ratio of 8.3.   If your company is expecting Earnings Per Share (EPS) of $3.43, what should the price if stock be?

Can you show me how to do this on a BA II Plus calculator?
Answer #1

1: Value of stock = D1/(Expected return - growth)

= 3.12/ (9%-3.1%)

=$52.88

2: First find the averge PE ratio

Enter X01= 7.3

X02= 8.6

X03= 9.1

X04= 7.1

X05= 5.9

X06= 8.3

Enter Xbar = 7.7167

Price of stock = PE Ratio *EPS

= 7.7167*3.43

= $ 26.47